# Coupled Requirements-driven Testing of CPS in Simulation and Reality Coupled Requirements-driven Testing of CPS: From Simulation To Reality

### v1

@@ V Research Roadmap
Model Homogeneity and Consistency: Aside from traceability, linking the different models is another crucial aspect, to ensure consistency among the requirements, the V&V, and the test model. This stems from the circumstance that different modeling formalisms are employed at the various stages, resulting in heterogeneous models, or even code snippets (e.g., when validating simulation results). Yet, the different models share identical properties and require trace links among them (cf. Traceability). If not properly managed, inconsistencies can be easily introduced throughout the whole process [38]. In traditional model-based development (MBD), commonly one modeling approach is used, however, the diverse nature of process levels and testing context (e.g., formal functional requirements testing vs. automated simulator testing vs. manual field testing) calls for support of a diverse set of models. As in the case of traceability, we leverage techniques from MBD, specifically ultra-larger scale heterogeneous model synchronization [39].
Representation of Test Stories: A structured and flexible description of different test stories is required. While model-based, and scenario-based testing are well-established techniques, solutions are typically limited to a narrow scope and usually target one execution environment; in our case, we have at least two - simulation and reality. Consequently, test stories need to cover a wide range from formal tests (cf. checking properties during simulation execution), to manually performing test steps at a field test with human interaction. For this purpose, we proposed the investigation of domain-specific languages (DSL) tailored towards test story specification, similar to the classic notion of a movie script which describes the environmental setting (cf. simulator or reality), the activities to happen (cf. instrumentation of the SuT), and eventually a test oracle for subsequent analysis. Capitalizing on DSLs substantially facilitates automation, allowing for intelligent test selection by covering a wide range of conditions with potentially limited resources (e.g., time-consuming simulation runs) [40].
@@
Another promising avenue of research in the context of Levels of Fidelity is to align them with already established measurement systems used to assess the maturity of a technology, e.g., the Technology readiness levels (TRLs) [42]. Incorporating such fine-grained levels into the safety analysis process, by making them part of the safety assurance case, would further strengthen the safety verdict depending on the reached level.
@@
Mapping Simulator Capabilities: Simulation environments currently cover a wide range of features and functionalities with various LoFs. While in the automotive domain, highly sophisticated, feature-rich simulation environments exist (e.g. CARLA), other areas still lack sophisticated simulation support. Particularly for sUAS systems, there is a need for standardized simulation support that enables methodological testing and validation [11]. To bridge the gap between simulation and testing, we are currently conducting an extensive literature analysis on sUAS simulation environments, with the goal of documenting commonalities and variability alongside the different features that are available, similar to the automotive domain [43]. This will not only deliver a better overview of existing solutions but — more importantly — provide the foundation for (automatically) mapping different LoFs. We envision a structured description of capabilities for simulation environments, similar to a feature model. For example, AirSim provides features for modeling and rendering 3D environments, including geospatial and weather data. This information can in turn be used to inform the LoF of a certain test story (or determine the desired LoF of a test story to be executed in a specific execution context), and automatically generate the respective setup and configuration scripts for test execution. This readily delivers a continuous testing process where simulation runs can be executed in a CI/CD DevOps process, to establish traceability among implemented and tested requirements, necessary simulator configurations, and corresponding test outcomes.
@@
Simulator-to-Reality Gap: Finally, the simulator-to-reality gap has long been recognized as a critical issue in robotics research [44]. It describes the mismatch of simulated and real physical-law performances caused by the inaccurate representation of the real environment in simulation [45]. In the event of our proposal, this issue becomes evidentily pertinent in moving testing from LoF-1 to LoF-3. Our approach however is not designed to reduce the simulator-to-reality gap but rather, as of its theoretical foundation, provide a tool to quantify this gap by correlating simulated and real execution traces. Put differently, our framework provides the measurement basis for aligning the fidelity of a simulation with the fidelity of reality, put simply, in how far results from simulation map to reality without loss of granularity.
@@ VI Conclusion and Future Work
Software Testing is a well-established research area, with mature processes, tools, and techniques. However, in sUAS and CPS domain, despite the existence of simulation environments commonly employed for testing, there is still a lack of support for executing a well-defined testing scenario, with an overall process for combining different types of tests, environmental factors, and properties. This often results in situations where systems are tested in inadequate simulation environments that do not adequately capture the real-world complexity, are only tested in a limited range of scenarios, and lack standard safety analysis processes. In this paper, we presented an initial framework for validating CPS, specifically focusing on sUAS and robotic applications, demonstrating its suitability by applying it to two different types of CPS. Based on our initial observations and insights we have compiled a research roadmap with five ongoing research activities related to model management and consistency, simulator capabilities, traceability, and structured test scenario representation.

### v2

@@ V Research Roadmap
Model Homogeneity and Consistency: Aside from traceability, linking the different models is crucial, to ensure consistency across requirements, the V&V, and the test model. This stems from the circumstance that different modeling formalisms are employed at the various stages, resulting in heterogeneous models, or even code snippets (e.g., when validating simulation results). Yet, the different models share identical properties and require trace links among them (cf. Traceability). If not properly managed, inconsistencies can be easily introduced throughout the whole process [40]. In traditional model-based development (MBD), commonly one modeling approach is used, however, the diverse nature of process levels and testing context (e.g., formal functional requirements testing vs. automated simulator testing vs. manual field testing) calls for support of a diverse set of models. As in the case of traceability, we leverage techniques from MBD, specifically ultra-larger scale heterogeneous model synchronization [41].
Representation of Test Stories: A structured and flexible description of different test stories is required. While model-based, and scenario-based testing are well-established techniques, solutions are typically limited to a narrow scope and usually target one execution environment; in our case, we have at least two - simulation and reality. Consequently, test stories need to cover a wide range from formal tests (cf. checking properties during simulation execution), to manually performing test steps at a field test with human interaction. For this purpose, we proposed the investigation of domain-specific languages (DSLs) tailored towards test story specification, similar to the classic notion of a movie script which describes the environmental setting (cf. simulator or reality), the activities to happen (cf. instrumentation of the SuT), and eventually a test oracle for subsequent analysis. Capitalizing on DSLs substantially facilitates automation, allowing for intelligent test selection by covering a wide range of conditions with potentially limited resources [42].
@@
Mapping Simulator Capabilities: Simulation environments currently cover a wide range of features and functionalities with various LoFs. While in the automotive domain, highly sophisticated, feature-rich simulation environments exist (e.g. CARLA), other areas still lack sophisticated simulation support. Particularly for sUAS, there is a need for standardized simulation support that enables methodological testing and validation [14]. To bridge the gap between simulation and testing, we are currently conducting an extensive literature analysis on sUAS simulation environments, with the goal of documenting commonalities and variability alongside the different features that are available, similar to the automotive domain [44]. This will not only deliver a better overview of existing solutions but — more importantly — provide the foundation for (automatically) mapping different LoFs. We envision a structured description of capabilities for simulation environments, similar to a feature model. For example, AirSim, based on Unreal Game Engine Technology, provides features for modeling and rendering 3D environments, including geospatial and weather data. This information can in turn be used to inform the LoF of a certain test story (or determine the desired LoF of a test story to be executed in a specific execution context), and automatically generate the respective setup and configuration scripts for test execution. As a result, simulation runs can be executed in a CI/CD DevOps process, allowing traceability between implemented requirements, simulator configurations, and test results.
@@
Simulator-to-Reality Gap: Finally, the simulator-to-reality gap has long been recognized as a critical issue in robotics research [45]. It describes the mismatch of simulated and real physical-law performances caused by the inaccurate representation of the real environment in simulation [46]. In our case, this issue becomes evidently pertinent in moving testing from LoF-1 to LoF-3. Our approach however is not designed to reduce the simulator-to-reality gap but rather, as of its theoretical foundation, provide a tool to quantify this gap by correlating simulated and real execution traces. Put differently, our framework provides the measurement basis for aligning the fidelity of a simulation with the fidelity of reality, put simply, in how far results from simulation map to reality without loss of granularity.
The above-mentioned six directions provide ample opportunities for improving the current state of the art in simulation-based validation of sUAS. While some research areas, like traceability, are well established, there is a need to explore automated approaches for supporting the traceability of simulation artifacts to assess and analyze sUAS safety. Additionally, integrating concepts from other domains, such as the proposed LoF simulator classification, requires an in-depth analysis of current simulator capabilities and a mapping to various types of domain- and system-specific requirements. Furthermore, our proposal is model-driven, allowing us to maintain consistency throughout the entire testing process, transferable across multiple domains, and providing opportunities to automate the process to the maximum extent possible.
@@ VI Conclusion and Future Work
Software Testing is a well-established research area, with mature processes, tools, and techniques. However, in the sUAS and CPS domain, despite the existence of simulation environments commonly employed for testing, there is still a lack of support for executing a well-defined testing scenario, with an overall process for combining different types of tests, environmental factors, and properties. This often results in situations where systems are tested in inadequate simulation environments that do not adequately capture the real-world complexity, are only tested in a limited range of scenarios, and lack standard safety analysis processes. In this paper, we presented an initial framework for validating CPS, specifically focusing on sUAS and robotic applications, demonstrating its suitability by applying it to two different types of CPS. Based on our initial observations and insights, we have compiled a research roadmap with five ongoing research activities related to model management and consistency, simulator capabilities, traceability, and structured test scenario representation.